OpenStudy (anonymous):

what is the least possible value of (x+2)^2 if -3 < or equal to x <0

OpenStudy (anonymous):

@calculusfunctions help me with this one now plz

OpenStudy (calculusfunctions):

Least possible value means minimum value. What is the minimum value of y = (x + 2)²?

OpenStudy (anonymous):

dy/dx=2(x+2)

OpenStudy (sirm3d):

the least value is 0, and this is when x = -2

OpenStudy (calculusfunctions):

Sure but you didn't need to do that! If y = a(x - h)² + k then the minimum value is k when x = h if a > 0 as @sirm3d correctly said but she should not give out answers!

OpenStudy (anonymous):

the picture is still not clear in my mind :(

OpenStudy (calculusfunctions):

If you understand the derivatives method better, then fine. Use that! So if dy/dx = 0, then x = ?

OpenStudy (anonymous):

x= -2

OpenStudy (calculusfunctions):

Yes! Now do you know how to use the first or second derivative tests to verify if this is a max or min?

OpenStudy (anonymous):

yes i do :)

OpenStudy (calculusfunctions):

OK then go ahead!